Lista dei presenti alla prova scritta del 10 settembre 2007.

Matricola  Votazione
10028447 Ammesso/a all’orale
10044524 Ammesso/a all’orale
10041531 Ammesso/a all’orale
10043090 Ammesso/a all’orale
10042734  Insufficiente